OpenStudy (rock_mit182):

hey guys please help me out with this question The electric field everywhere on the surface of a thin spherical shell of radius 0.750 m is measured to be equal to 890 N/C and points radially toward the center of the sphere. (a) What is the net charge within the sphere’s surface? (b) What can you conclude about the nature

OpenStudy (rock_mit182):

@Luigi0210 @Nnesha

OpenStudy (rock_mit182):

@sissyedgar @YanaSidlinskiy

OpenStudy (sissyedgar):

idk!

OpenStudy (rock_mit182):

NO prob ! ;) i need some help please someone help me out

OpenStudy (rock_mit182):

@M4thM1nd

OpenStudy (anonymous):

Ok, first of all, let's use the Gauss law: \[EA = Q/\epsilon_0\], but remember that it is a spherical shell, not a sphere, that means that the inner portion is empty.

OpenStudy (anonymous):

First let's use Gauss law for \(r \ge R\), where \(R\) is the spherical shell radius

OpenStudy (anonymous):

\(E*(4\pi r^2)=Q/\epsilon_0\), so... \[E = \frac{Q}{4\pi \epsilon_0 r^2}\] For \(r \ge R\)

OpenStudy (anonymous):

Now, for the inner part, we have no charges enclosed by a gaussian surface, so.. \(EA = 0 \Rightarrow E = 0\), for \(r < R\)

OpenStudy (anonymous):

As the electric field points towards teh center, the charge on the surface is Negative, using the fist case \(r \ge R\), we get: \[Q = 4 \pi \epsilon_0 R^2E = 4 \pi \epsilon_0* 0.750^2*(-890) =...\]

OpenStudy (anonymous):

remember that \(\epsilon_0 = 8.85\times 10^{-12} C^2/{N\cdot m^2}\)

OpenStudy (anonymous):

Got it @rock_mit182 ?

OpenStudy (rock_mit182):

@M4thM1nd what if there is a charge, in the center of an sphere, with an electric field pointing inwards, How would I calculate the net charge (q NET) ?

OpenStudy (anonymous):

Can you be more specific? Is it a pontual charge in a spherical shell?

OpenStudy (anonymous):

if so, again we use Gauss Law, \[EA = Q_{env}/\epsilon_0\] \[E*(4\pi r^2) = Q_{env}/\epsilon_0\] \[E = \frac{Q_{env}}{4\pi \epsilon_0 r^2}\], for \(r> 0\)

OpenStudy (rock_mit182):

thanks for helping

OpenStudy (anonymous):

No problem ;)